# TU Munich, Germany

### v1

@@ Abstract
Precise anomaly detection in medical images is critical for clinical decision-making. While recent unsupervised or semi-supervised anomaly detection methods trained on large-scale normal data show promising results, they lack fine-grained differentiation, such as benign vs. malignant tumors. Additionally, ultrasound (US) imaging is highly sensitive to devices and acquisition parameter variations, creating significant domain gaps in the resulting US images. To address these challenges, we propose UltraAD, a vision-language model (VLM)-based approach that leverages few-shot US examples for generalized anomaly localization and fine-grained classification. To enhance localization performance, the image-level token of query visual prototypes is first fused with learnable text embeddings. This image-informed prompt feature is then further integrated with patch-level tokens, refining local representations for improved accuracy. For fine-grained classification, a memory bank is constructed from few-shot image samples and corresponding text descriptions that capture anatomical and abnormality-specific features. During training, the stored text embeddings remain frozen, while image features are adapted to better align with medical data. UltraAD has been extensively evaluated on three breast US datasets, outperforming state-of-the-art methods in both lesion localization and fine-grained medical classification. The code will be released upon acceptance.
@@ 1 Introduction
Medical ultrasound (US) is a widely used imaging modality for examining internal organs, such as the breast and thyroid, due to its real-time capability, non-radiative nature, and accessibility. In remote or low-income regions, it is often the only available diagnostic tool. However, US images frequently suffer from low quality and significant domain variations due to differences in devices and imaging parameters [17, 5]. To address these challenges, a generalized anomaly detection (AD) algorithm that adapts across diverse anatomical structures and imaging domains is highly demanded for supporting clinical decision-making, particularly for junior clinicians with limited clinical experience.
While supervised deep learning has achieved phenomenal success in medical image segmentation [12, 18, 33, 23, 11, 15], these methods require large labeled training data, which is costly in the US imaging field. To overcome this limitation, unsupervised and self-supervised approaches leveraging autoencoders and generative models have been developed using only unlabeled data [4, 3, 2]. Despite success in MRI anomaly detection, these methods cannot be directly adopted to US images because of the noisy nature and limited imaging field of view.
@@
In this study, we introduce UltraAD, a CLIP-based framework for lesion localization and fine-grained anomaly classification through few-shot adaptation. Inspired by multi-task learning in medical image segmentation [26, 16], UltraAD unifies pixel-wise anomaly localization and image-level classification to boost performance across tasks. To address the domain gap between natural and US images, a small set of US samples (4/8/16 shots in this study) is needed to adapt the CLIP model. To the best of our knowledge, this is the first CLIP-based model tailored for meeting US diagnosis requirements, leveraging domain-specific pathological information to improve generalization across imaging variations. Notably, the few-shot adaptation is performed on a single public breast dataset, while validation is conducted on two unseen breast datasets acquired using different machines, patients, and probe types. This highlights the robustness and effectiveness of the proposed method in real scenarios. The key contributions are: (1) A simple yet effective few-shot adaptation approach that generalizes across unseen US images of the same anatomy, despite significant domain variations in machine types and acquisition parameters; (2) a vision-language fusion strategy, combining learnable prompt tokens and image embeddings at both global (image-level) and local (patch-wise) scales to enhance anomaly localization; (3) a feature memory bank built from a small set of paired US images and text descriptions, enabling enhanced anomaly classification
@@ 2 Methodology
Figure 1: Proposed pipeline. Via pretrained image and text encoders, text token embeddings $P_{t}$, image token embeddings $F_{t}$ and onehot label $L_{t}$ are preloaded from the S-shot C-class dataset. A mini-net projects image token embeddings $f$ to obtain $f^{\prime}$. For anomaly classification, $F_{t}$, $P_{t}$ and $L_{t}$ are utilized to generate a classification score with $f^{\prime}$ through feature similarity. For anomaly localization, two initial text prompts indicating normal and abnormal are added with the projected image token embedding $f^{\prime}$. Two localization masks are produced by calculating the cosine similarity between an updated image token embedding $w^{\prime}$ and the projected patch token $f^{\prime}_{p}$, and a text-image fused text-embedding $w_{fuse}$ with the original patch token $f_{p}$. The final anomaly map prediction is computed as $\hat{Y}=\frac{1}{2}(\hat{Y}_{1}+\hat{Y}_{1})$
@@ 2.1 CLIP for Anomaly Detection
[…]
$\hat{y}=\frac{\exp(\langle\mathbf{w}_{a},f\rangle)}{\exp(\langle\mathbf{w}_{n},f\rangle)+\exp(\langle\mathbf{w}_{a},f\rangle)}~{},\quad\hat{Y}=\sigma(\frac{\exp(\langle\mathbf{w}_{a},f_{p}\rangle)}{\exp(\langle\mathbf{w}_{n},f_{p}\rangle)+\exp(\langle\mathbf{w}_{a},f_{p}\rangle)})$
@@
$P=[\text{A}][\text{photo}][\text{of}][\text{a}][\textit{state}][u+f^{\prime}]$
@@ 2.2 Memory-Boosted Few-Shot Adaptation
$\hat{y}=f^{\prime}\textbf{P}_{\text{t}}^{T}+\exp(f^{\prime}\mathbf{F}_{\text{t}}^{T}-1)L_{\text{t}},$ (2)
@@ 2.4 Multitask Learning for Anomaly Classification and Localization
The class token, which acts as an embedding containing global image information, can be used for image classification and integrated into the prompt for conditioned prompt learning. To achieve this, we combine anomaly classification and anomaly localization by using a shared, dynamically updated class token within a unified framework. This approach seamlessly integrates anomaly classification with anomaly localization within a cohesive learning pipeline. The initially extracted CLIP image feature, denoted as $f$, is projected via a MiniNet $\Phi(\cdot)$, resulting in $f^{\prime}$. During training, UltraAD refines the pixel-level anomaly maps $\hat{Y}_{1}$ and $\hat{Y}_{2}$ using a combination of Dice loss [20] and Focal loss [19], equally weighted, on auxiliary data (few-shot samples) to ensure accurate segmentation. Meanwhile, it refines classification performance by employing cross-entropy loss. The image feature memory $\textbf{F}_{t}$ will be refined during the training while text feature $\textbf{P}_{t}$ remains frozen. The shared feature $f^{\prime}$ is jointly refined, which will be used to compute the classification result as well as the segmentation result.
@@ 3.1 Experimental Setup
Following the evaluation methods used in anomaly detection and localization tasks [14, 32, 6, 21], we utilize the AUROC score for both anomaly classification and localization. All the values in Tab.3.1 and Tab.2 shows the average I-AUROC/P-AUROC of three experiments using 3 seeds for few-shot sample selection to ensure fair evaluation and reduce random few-shot sample variations.
Anomaly Detection (binary) Anomaly Detection (binary)
@@
3.2 Performance Analysis
Anomaly Detection and Classification We conduct comprehensive evaluations on two distinct few-shot classification tasks: anomaly detection (binary) and anomaly classification (multi-class). For anomaly detection, we formulate the task as a binary classification problem and evaluate against leading CLIP-based AD approaches [14, 32, 6, 21, 10]. Tab.3.1 (left) demonstrate UltraAD’s superior performance across all settings, particularly in low-shot scenarios where it achieves a substantial 10% gain over baselines in the challenging 4-shot setting. In addition to anomaly detection, we address the more complex task of anomaly classification, which involves fine-grained classification of anomaly types under limited data constraints. Comparing against state-of-the-art CLIP-based few-shot learning techniques including advanced prompt tuning and adapter-based methods[31, 13, 9, 28], our results in Tab.3.1 (right) shows UltraAD’s significant advantages, especially on the complex BUSZS dataset, achieving 80% score on most settings. We attribute this superior performance to our novel integration of segmentation-guided learning. Additionally, we employ a mask-guided post-processing technique, commonly used in recent anomaly detection methods [14, 32], which utilizes segmentation masks to refine the separation between normal and abnormal classes. This approach further enhances overall performance, specifically using the formulation $y_{pred}=\frac{1}{2}(\max(\hat{Y})+\hat{y})$.
@@ 3.2 Performance Analysis
Anomaly Localization Tab. 2 presents a comprehensive analysis of our method against state-of-the-art baseline models for anomaly localization. Our proposed UltraAD demonstrates superior performance compared to most existing methods. AnomalyCLIP[32] demonstrates competitive performance attributed to its V-V attention mechanism, which significantly enhances local feature perception. Nevertheless, AnomalyCLIP exhibits limitations in image level anomaly detection show in Tab.3.1. Fig.2 presents the quantitative results of localization. WinCLIP[14] and VCPCLIP[21] failed to produce satisfactory maps, while MVFA[10] and AnomalyCLIP[32] struggled to generate continuous regions as anomaly areas. AdaCLIP[6] included some normal regions in its predictions. Our method detects anomaly locations and predicts them as continuous regions, which can be further used for precise segmentation.
@@ 3.3 Ablation Study
We evaluate UltraAD modules via an ablation study in a 4-shot setting (M3) on the BUSZS. The results are reported using the notation (I-AUROC/P-AUROC). Our final method achieves scores of (82.3/93.4). Specifically, we examine the impact of employing US-specific terminologies by using US-unaware prompts to generate $\textbf{P}_{t}$, resulting in scores of (72.3/88.6). To verify the functionality of Memory-Boosted Few-Shot Adaptation, we remove classification and focus on segmentation loss, reducing localization performance and yielding (-/84.3) without joint learning. To verify the Patch-Wise Image-Language Fusion, we omit the Patch-Wise Image-Language Fusion module, instead directly utilizing the linearly projected patch tokens $f^{\prime}_{p}$ for anomaly localization, resulting in (71.2/90.3).

### v2

@@ Abstract
Precise anomaly detection in medical images is critical for clinical decision-making. While recent unsupervised or semi-supervised anomaly detection methods trained on large-scale normal data show promising results, they lack fine-grained differentiation, such as benign vs. malignant tumors. Additionally, ultrasound (US) imaging is highly sensitive to devices and acquisition parameter variations, creating significant domain gaps in the resulting US images. To address these challenges, we propose UltraAD, a vision-language model (VLM)-based approach that leverages few-shot US examples for generalized anomaly localization and fine-grained classification. To enhance localization performance, the image-level token of query visual prototypes is first fused with learnable text embeddings. This image-informed prompt feature is then further integrated with patch-level tokens, refining local representations for improved accuracy. For fine-grained classification, a memory bank is constructed from few-shot image samples and corresponding text descriptions that capture anatomical and abnormality-specific features. During training, the stored text embeddings remain frozen, while image features are adapted to better align with medical data. UltraAD has been extensively evaluated on three breast US datasets, outperforming state-of-the-art methods in both lesion localization and fine-grained medical classification. Project page: https://karolinezhy.github.io/UltraAD/
@@ 1 Introduction
Medical ultrasound (US) is a widely used imaging modality for examining internal organs, such as the breast and thyroid, due to its real-time capability, non-radiative nature, and accessibility. In remote or low-income regions, it is often the only available diagnostic tool. However, ultrasound images frequently suffer from low quality and significant domain variations due to differences in imaging devices and acquisition parameters [16, 4], which pose challenges for ultrasound image understanding [18]. To address these challenges, a generalized anomaly detection (AD) algorithm that adapts across diverse anatomical structures and imaging domains is highly demanded for supporting clinical decision-making, particularly for junior clinicians with limited clinical experience.
While supervised deep learning has achieved phenomenal success in medical image segmentation [11, 17, 33, 23, 10, 14], these methods require large labeled training data, which is costly in the US imaging field. To overcome this limitation, unsupervised and self-supervised approaches leveraging autoencoders and generative models have been developed using only unlabeled data [3, 2]. Despite success in MRI anomaly detection, these methods cannot be directly adopted to US images because of the noisy nature and limited imaging field of view.
@@
In this study, we introduce UltraAD, a CLIP-based framework for lesion localization and fine-grained anomaly classification through few-shot adaptation. Inspired by multi-task learning in medical image segmentation [26, 15], UltraAD unifies pixel-wise anomaly localization and image-level classification to boost performance across tasks. To address the domain gap between natural and US images, a small set of US samples (4/8/16 shots in this study) is needed to adapt the CLIP model. To the best of our knowledge, this is the first CLIP-based model tailored for meeting US diagnosis requirements, leveraging domain-specific pathological information to improve generalization across imaging variations. Notably, the few-shot adaptation is performed on a single public breast dataset, while validation is conducted on two unseen breast datasets acquired using different machines, patients, and probe types. This highlights the robustness and effectiveness of the proposed method in real scenarios. The key contributions are: (1) A simple yet effective few-shot adaptation approach that generalizes across unseen US images of the same anatomy, despite significant domain variations in machine types and acquisition parameters; (2) a vision-language fusion strategy, combining learnable prompt tokens and image embeddings at both global (image-level) and local (patch-wise) scales to enhance anomaly localization; (3) a feature memory bank built from a small set of paired US images and text descriptions, enabling enhanced anomaly classification.
@@ 2 Methodology
Figure 1: Proposed pipeline. Via pretrained image and text encoders, text token embeddings $P_{t}$, image token embeddings $F_{t}$ and onehot label $L_{t}$ are preloaded from the S-shot C-class dataset. A mini-net projects image token embeddings $f$ to obtain $f^{\prime}$. For anomaly classification, $F_{t}$, $P_{t}$ and $L_{t}$ are utilized to generate a classification score with $f^{\prime}$ through feature similarity. For anomaly localization, two initial text prompts indicating normal and abnormal are added with the projected image token embedding $f^{\prime}$. Two localization masks are produced by calculating the cosine similarity between an updated image token embedding $w^{\prime}$ and the projected patch token $f^{\prime}_{p}$, and a text-image fused text-embedding $w_{fuse}$ with the original patch token $f_{p}$. The final anomaly map prediction is computed as $\hat{Y}=\frac{1}{2}(\hat{Y}_{1}+\hat{Y}_{1}).$
@@ 2.1 CLIP for Anomaly Detection
To adapt CLIP [22] for anomaly detection, two contrasting text prompts are used: A photo of a normal [CLS] and A photo of an abnormal [CLS], where [CLS] represents the specific image class category, e.g., breast US image. The text encoder $\mathit{E_{T}}(\cdot)$ processes text prompts to generate the text embedding $\mathbf{w}_{n}\in\mathbb{R}^{D}$ for the normal class and $\mathbf{w}_{a}\in\mathbb{R}^{D}$ for the abnormal class. Given an input image $X$, the image encoder $\mathit{E_{I}}(\cdot)$ outputs a global class token $f\in\mathbb{R}^{D}$ and local patch tokens $f_{p}\in\mathbb{R}^{h\times w\times D}$, where $h$ and $w$ denote the spatial dimensions in feature space. The anomaly scores are obtained based on cosine similarity $\langle\cdot\rangle$ between text and visual embedding for both global image-level and local pixel-level, i.e.,
$\hat{y}=\frac{\exp(\langle\mathbf{w}_{a},f\rangle)}{\exp(\langle\mathbf{w}_{n},f\rangle)+\exp(\langle\mathbf{w}_{a},f\rangle)},\hat{Y}=\sigma(\frac{\exp(\langle\mathbf{w}_{a},f_{p}\rangle)}{\exp(\langle\mathbf{w}_{n},f_{p}\rangle)+\exp(\langle\mathbf{w}_{a},f_{p}\rangle)})$ (1)
@@
$P=[\text{A}][\text{photo}][\text{of}][\text{a/an}][\textit{state}][u+f^{\prime}]$
@@ 2.2 Memory-Boosted Few-Shot Adaptation
$\hat{y}=f^{\prime}\textbf{P}_{\text{t}}^{T}+\exp(f^{\prime}\mathbf{F}_{\text{t}}^{T}-1)L_{\text{t}}$ (3)
@@ 2.4 Multitask Learning for Anomaly Classification and Localization
The class token, which acts as an embedding containing global image information, can be used for image classification and integrated into the prompt for conditioned prompt learning. To achieve this, we combine anomaly classification and anomaly localization by using a shared, dynamically updated class token within a unified framework. This approach seamlessly integrates anomaly classification with anomaly localization within a cohesive learning pipeline. The initially extracted CLIP image feature, denoted as $f$, is projected via a MiniNet $\Phi(\cdot)$, resulting in $f^{\prime}$. During training, UltraAD refines the pixel-level anomaly maps $\hat{Y}_{1}$ and $\hat{Y}_{2}$ using a combination of Dice loss [20] and Focal loss [19], equally weighted, on auxiliary data (few-shot samples) to ensure accurate segmentation. Meanwhile, it refines classification performance by employing cross-entropy loss. The image feature memory $\textbf{F}_{t}$ will be refined during the training while text feature $\textbf{P}_{t}$ remains frozen. The network is trained end-to-end with a joint optimization of segmentation and classification losses. The shared feature $f^{\prime}$ is jointly refined, which will be used to compute the classification result as well as the segmentation result.
@@ 3.1 Experimental Setup
Following the evaluation methods used in anomaly detection and localization tasks [13, 32, 5, 21], we utilize the AUROC score for classification and AUROC/AUPRC for localization. All the values in Tab. 2 and Tab. 3 shows the average metrics of three experiments using 3 seeds for few-shot sample selection to ensure fair evaluation and reduce random few-shot sample variations.
3.2 Performance Analysis
Table 2: Comparison of anomaly detection (left) and classification (right) performance across benchmarks, evaluated using image-level AUROC score. The best values are bold, and the second best is underlined.
Anomaly Detection (binary) Anomaly Detection (multi-class)
@@ 3.2 Performance Analysis
[…]
@@
Anomaly Localization Tab. 3 presents a comprehensive analysis of our method against state-of-the-art baseline models for anomaly localization. Our proposed UltraAD demonstrates superior performance compared to most existing methods. AnomalyCLIP[32] demonstrates competitive performance attributed to its V-V attention mechanism, which significantly enhances local feature perception. Nevertheless, AnomalyCLIP exhibits limitations in image level anomaly detection show in Tab.2. Fig.2 presents the quantitative results of localization. WinCLIP[13] and VCPCLIP[21] failed to produce satisfactory maps, while MVFA[9] and AnomalyCLIP[32] struggled to generate continuous regions as anomaly areas. AdaCLIP[5] included some normal regions in its predictions. Our method detects anomaly locations and predicts them as continuous regions, which can be further used for precise segmentation.
@@ 3.3 Ablation Study
We evaluate UltraAD modules via an ablation study in a 4-shot (one-seed) setting on BUSZS Dataset. The results are reported using the notation (I-AUROC/P-AUROC). Our final method achieves scores of (82.3/93.4). Specifically, we examine the impact of employing US-specific terminologies by using US-unaware prompts to generate $\textbf{P}_{t}$, resulting in scores of (72.3/88.6). To verify the functionality of Memory-Boosted Few-Shot Adaptation, we remove classification and focus on segmentation loss, reducing localization performance and yielding (-/84.3) without joint learning. To verify the Patch-Wise Image-Language Fusion, we omit the Patch-Wise Image-Language Fusion module, instead directly utilizing the linearly projected patch tokens $f^{\prime}_{p}$ for anomaly localization, resulting in (71.2/90.3).